OpenStudy (anonymous):

If a book has 252 pages how many digits have been used to number the pages?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

From Page Number 1 to 9 we use 1 digit per page => digit used = 9 From page Number 10 to 99 we use 2 digit per page => digit used = 90 *2 = 180 From page Number 100 to 252 we use 3 digit per page => digit used = 153 *3 = 459 Therefore total digit used = 9+180+459 = 648

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Sorry, the digits 0 to 9 have been used to number the pages.. Your question said nothing about having to count them more than once.

OpenStudy (anonymous):

People who gave u a medal for wrongly answering ur own question...duh!

OpenStudy (anonymous):

How come 0 is used to number the page ???? I haven't seen any book in which zero is used as the first page.

OpenStudy (anonymous):

204.

OpenStudy (anonymous):

Nothing to say?

OpenStudy (anonymous):

204